A sample of HCI gas is placed in an electric field of 2.5 × 104 N/C. The dipole moment of each HCI molecule is 3.4 × 10-30 C-m. Find the maximum torque that can act on a molecule.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. E = 2.5 × 104 P = 3.4 × 10^–30 τ = PE sin θ = P × E × 1 = 3.4 × 10^–30 × 2.5 × 10^4 = 8.5 × 10^–26